OpenStudy (anonymous):

For the polynomial function h(x) = 2x^4 - x^3 + 2x^2 - 4x - 3, evaluate the following: h(3)

OpenStudy (anonymous):

h(3) = 2*81 - 27 + 2*9 - 4*3 -3 = 162 - 27 + 18 -12 -3 = 180 - 42 = 138